OpenStudy (anonymous):

Give an example: A sequence Sn of positive numbers with no upper bound which does not diverge to infinity.

OpenStudy (anonymous):

Does this work: (-1)^(2n) ____________ n

OpenStudy (loser66):

nope, (-1)^2n =1, the sequence has 1 is an upper bound

OpenStudy (anonymous):

1/n ?

OpenStudy (anonymous):

or is 1 still an upper bound in this case?

OpenStudy (loser66):

\(\dfrac{(-1)^{2n}}{n}=\dfrac{1}{n}\) has 1 is an upperbound